### v1

# From Templates to Natural Language: Generalization Challenges in Instruction-Tuned LLMs for Spatial Reasoning

Abstract
Instruction-tuned large language models (LLMs) have shown strong performance on a variety of tasks; however, generalizing from synthetic to human-authored instructions in grounded environments remains a challenge for them. In this work, we study generalization challenges in spatial grounding tasks where models interpret and translate instructions for building object arrangements on a $2.5$D grid. We fine-tune LLMs using only synthetic instructions and evaluate their performance on a benchmark dataset containing both synthetic and human-written instructions. Our results reveal that while models generalize well on simple tasks, their performance degrades significantly on more complex tasks. We present a detailed error analysis of the gaps in instruction generalization.
From Templates to Natural Language: Generalization Challenges in Instruction-Tuned LLMs for Spatial Reasoning
Chalamalasetti Kranti1, Sherzod Hakimov1, David Schlangen1,2 1Computational Linguistics, Department of Linguistics University of Potsdam, Germany 2German Research Center for Artificial Intelligence (DFKI), Berlin, Germany {kranti.chalamalasetti, sherzod.hakimov, david.schlangen}@uni-potsdam.de
1 Introduction
In recent years, LLMs have demonstrated impressive capabilities in following natural language instructions across a wide range of tasks (Brown et al., 2020; Yang et al., 2024; Firoozi et al., 2025). However, their ability to generalize from training data to human-authored instructions remains limited in grounded environments, such as collaborative block manipulation (Volovikova et al., 2024), table setting (Shridhar et al., 2020), and structure building (Chalamalasetti et al., 2024; Chaturvedi et al., 2024), where language must be translated into precise spatial actions.
In contrast, studies in human communication (Fischer and Moratz, 2001; Li et al., 2016; Marge et al., 2017) show that people adapt their language depending on the context and the intended partner. Speakers tend to provide step-by-step directions and fewer perspective-dependent expressions when instructing a robot.
Yet despite the simplified instructions, instruction-tuned LLMs often struggle (Kranti et al., 2024) with spatial tasks when evaluated on human-authored language. This raises a question: if human speakers adapt their language to be more robot-friendly, why do models trained on such data still fail to generalize?
Figure 1: Fine-tuned LLMs trained on synthetic instructions are evaluated on both synthetic (SI) and human (HI) instructions. A virtual simulator verifies whether generated response matches the gold configuration.
For instance, the same speaker might say “put the red block on top of the blue one” in one situation, and “make a staircase” in another. These variations challenge model generalization across the linguistic spectrum.
We investigate this question in the context of structure-building tasks, where instructions guide the manipulation. To better understand the limits of instruction generalization, we fine-tune LLMs (see Figure 1) using synthetic, template-based instructions that reflect the explicit and unambiguous style. We examine whether synthetic-only fine-tuning can enable robust generalization by comparing model performance across tasks that vary in linguistic complexity and spatial structure.
We leverage an existing dataset (Kranti et al., 2024) that provides both synthetic and human-authored instructions for the same grounded spatial reasoning task. Each task involves building specific arrangements of colored shapes on a 2.5D grid (8x8), referred to as “boards.” Every board is annotated with synthetic and human-authored instructions, enabling direct comparison of model behavior across instruction styles in a shared task.
We fine-tune instruction-tuned LLMs using only the synthetic instructions and evaluate their ability to generalize to human-authored ones. Our results show that while models perform well on simple boards, achieving high execution scores, performance drops significantly on regular boards. This suggests that instruction generalization is closely tied to the compositional and structural demands of the underlying task and that fine-tuning only on synthetic datasets is insufficient for robust transfer to more complex spatial concepts.
Our contributions are: (a) we conduct a systematic analysis of generalization to human instructions in structure-building tasks; (b) we show that LLMs fine-tuned on synthetic data perform poorly on human-authored instructions for repetitive tasks.
2 Related Work
Synthetic vs. Natural Language Generalization: instruction-tuned models often struggle to generalize from synthetic data to natural language (Li et al., 2023; Nwankwo et al., 2025; Shi et al., 2025; Nadas et al., 2025). This highlights the challenges of linguistic complexity in real-world instructions. Our work extends this to the spatial grounding task, showing similar failures in generalization despite high performance on synthetic tasks.
Fine-tuning on Synthetic and Human Instructions in Structure-Building Tasks: prior work has explored fine-tuning LLMs for collaborative structure building in environments like Minecraft (Chalamalasetti et al., 2024; Chaturvedi et al., 2024). These studies typically fine-tune on a mix of synthetic and human-authored instructions but still report limited generalization. Furthermore, Minecraft dataset does not distinguish based on complexity and structures with and without object repetitions, making it difficult to isolate the impact of fine-tuning. Our work builds on these by analyzing performance gaps across instruction styles and board types that affect generalization.
Model In-domain Out-of-domain
Simple Boards Regular Boards Seen Unseen
Before After Before After Before After Before After
CodeLLama-7B 0.09 0.32 0.02 0.05 0.56 0.03 0.55 0.04
CodeLlama-34B 0.10 0.26 0.02 0.12 0.16 0.06 0.17 0.07
Qwen2.5-Coder-7B 0.07 0.59 0.00 0.25 0.60 0.00 0.55 0.00
Qwen2.5-Coder-32B 0.14 0.48 0.20 0.23 0.86 0.94 0.82 0.95
Qwen3-8B 0.00 0.54 0.00 0.11 0.00 0.00 0.00 0.00
Qwen3-32B 0.00 0.48 0.00 0.08 0.00 0.00 0.00 0.00
Llama3.3-70B 0.11 0.78 0.07 0.15 0.00 0.16 0.00 0.20
Claude-3 0.43 - 0.25 - - - - -
GPT-4o 0.17 - 0.31 - - - - -
Human baseline 0.98 - 0.76 - - - - 0.95
Table 1: Execution success rates before and after fine-tuning. In-domain: human instructions from the training domain; Out-of-domain: TidyBot dataset: Seen and Unseen sets. Best post-fine-tuning scores are shown in blue.
3 Task Formulation and Datasets
We focus on a spatial reasoning task where the objective is to execute a sequence of spatial actions based on natural language instruction. Our primary dataset is from (Kranti et al., 2024), which involves building a board i.e., arranging objects on a grid. An object is a structured combination of shapes (e.g., washers, screws, nuts, and bridges) arranged in a specific pattern. Each board is categorized by its type: either simple (non-repetitive arrangements) or regular (repetitive arrangements).
Figure 2: An example task illustrating the alignment between the code representation (left), its target structure (top-right), and two instruction styles (bottom).
Each task instance (see Figure 2) consists of an instruction paired with a target code representation, a Python function describing the intended final arrangement on the grid. The goal of the model is to generate this function based on the given instruction and environment context. Furthermore, each task instance is associated with two variants of instructions: (a) synthetic instructions, automatically generated using templated grammar. These are structurally consistent, semantically unambiguous, and direct; (b) human-written instructions collected from annotators (see Appendix A). These use free-form natural language to describe the same goals as the synthetic instructions.
We prompt instruction-tuned language models with both instruction types and evaluate the resulting code by executing it in a virtual simulated environment. The generated board is then compared to the ground truth to compute execution success—the proportion of cases where the model’s output yields the correct spatial configuration.
To test generalization, we fine-tune models on the training set of synthetic instructions and evaluate performance on both synthetic and human-written instructions. This setup allows us to study how well instruction-tuned LLMs generalize from synthetic (templated) instructions to natural human language within the same domain and across tasks. To assess cross-domain generalization, we evaluate these models on the TidyBot (Wu et al., 2023) dataset, which comprises natural language instructions for tidying up real-world household environments. The instances in this dataset are not seen during training (both subsets: seen & unseen).
4 Experimental Setup
We evaluate a diverse set of pre-trained LLMs with varying sizes and architectures, focusing on their ability to interpret grounded spatial instructions and generate executable action sequences. These models are used as-is and fine-tuned on synthetic data only, without exposure to human-written instructions during training.
Models Evaluated
We include both code-centric models (CodeLlama-7B, 34B; Qwen2.5-Coder-7B, 32B) and general-purpose models (Qwen3-8B, 32B; LLaMA-3.3-70B). All models are loaded in 4-bit precision using the Unsloth library [^1] https://unsloth.ai/. We compared these results against the reported scores of the closed models GPT-4o and Claude-3 (version-opus) by Kranti et al. (2024).
Fine-Tuning
Models are fine-tuned on synthetic instructions paired with target code representations. The training set comprises a mixture of Simple Boards (1072 boards) and Regular Boards (1168 boards) (Kranti et al., 2024). We adopt a chat-style prompt format, including environment details (more details in Appendix A.1), followed by the instruction and target code. We use the QLoRA configuration provided by Unsloth with the following hyperparameters: 1 epoch, 2 steps, adam-8bit optimizer, rank $r=16$, lora_alpha=16, lora_dropout=0.0, batch size=2, learning rate=$2\mathrm{e}{-4}$.
Evaluation Metrics
We assess model performance on instruction-following capability, execution accuracy, and cross-domain robustness. Our goal is to evaluate whether models can generalize across instruction types (synthetic $\rightarrow$ human) and across domains (arranging board $\rightarrow$ sorting objects).
Error Rate: A metric capturing instruction adherence based on predefined response constraints (output formatting) in the prompt (Appendix A.1).
Execution Success Rate: Our primary task success metric. A prediction is successful if the board configuration in the generated response matches the gold sequence when executed in a deterministic simulator. This involves matching correct component usage, spatial placement, and ordering.
Figure 3: Error analysis on Simple Boards (top) and Regular Boards (bottom), with outputs from specific models indicated above each result. Errors highlight challenges in spatial generalization from natural language instructions.
5 Results
The results are presented in Table 1. We present results for (a) generalization from synthetic to human instructions on the structure-building dataset and (b) transfer to human-written instructions in the TidyBot dataset (Wu et al., 2023).
5.1 In-Domain Generalization: Synthetic → Human
Table 1 reports execution success, and Table 5 (in Appendix A.2.3) reports error rates on simple and regular boards for each model for human instructions and the findings are: (a) fine-tuning substantially improves success on human instructions, particularly for models like Qwen2.5-7B, Qwen2.5-32B, and LLaMA-3.3-70B; (b) models perform better on simple boards than on regular boards, indicating that object repetition (and resulting referential ambiguity) challenges generalization; (c) large models such as GPT-4 and Claude-3, evaluated few-shot, show moderate ability to generalize to human instructions but underperform compared to fine-tuned models. High human baseline scores indicate that the task has a clear upper bound, which current models have yet to reach.
5.2 Cross-Domain Generalization: Transfer to TidyBot
To test whether models generalize beyond vocabularies seen during training, we evaluate models fine-tuned on synthetic structured boards on another domain dataset: the TidyBot dataset (Wu et al., 2023), which contains free-form human instructions for real-world object arrangement. The results are reported in Table 1 and the findings are: (a) top-performing models (Qwen2.5-32B, LLaMA-3.3-70B) maintain high success show generalization to unseen instructions and unseen objects, suggesting that fine-tuning on well-structured, synthetic spatial tasks can yield transferable skills; (b) other models (e.g., Qwen2.5-7B, CodeLlama-7B) degrade significantly, indicating possible overfitting.
Fine-tuning on synthetic instructions enables strong generalization to human-written instructions in low-ambiguity environments. Referential ambiguity (see Appendix A.3), introduced by repetitive objects, significantly correlates with failures. Some models learn instruction-to-structure mappings that generalize across language form (synthetic → human) and object domain (seen → unseen).
5.3 Qualitative Error Analysis
We categorize incorrect executions across models by error type (see Figure 3). The analysis indicates: (a) Element mismatch is most common on synthetic instructions; (b) Hierarchical instructions and ambiguous references dominate human instructions in regular boards, reinforcing the idea that repetition introduces referential entropy; (c) Models like Llama-3.3-70B and Qwen2.5-32B exhibit fewer invalid sequences, suggesting that fine-tuning improves overall generalization.
6 Conclusion
This paper investigates when and how synthetic instruction data can support generalization to human-authored instructions in grounded spatial reasoning tasks. Our findings show that synthetic-only fine-tuning can enable generalization when the arrangement does not involve object repetitions. Through semantic similarity analysis and error categorization, we identified referential ambiguity as a bottleneck: performance declines when instructions contain linguistic variations due to object repetitions. Future work should focus on developing synthetic datasets that mimic these linguistic variations.
Limitations
Our study focuses on generalization from synthetic to human instructions in spatial reasoning tasks, but several limitations remain. First, the training data is entirely synthetic and rule-based; while it enables controlled supervision, it lacks the linguistic diversity, ambiguity, and noise characteristic of real-world language. Second, the target code representation is highly task-specific, designed for grid-based pick-and-place operations, which may limit transferability to other spatial reasoning domains or instruction-following tasks with different semantics. Third, although our evaluation includes human-authored instructions, they are limited to single-turn settings and do not capture the challenges of multi-turn or collaborative spatial interactions.
Appendix A Appendix
A.1 Prompt Templates
In our task setup, LLMs are prompted to translate natural language instructions into structured code representations. We adopt a chat-style format to align with the pretraining of instruction-tuned models and design a multi-part prompt (see Figure 7 and Figure 8) that includes details of the $2.5$D grid environment along with available API functions. The prompt structure differs slightly between Simple and Regular boards—particularly in terms of response constraints and the nature of in-context examples provided. This same prompt format is used both during inference and fine-tuning.
A.2 Fine-Tuning
Our dataset consists of spatial instruction-target code pairs over two environment types: Simple Boards (SB) and Regular Boards (RB). Table 2 summarizes the distribution across training, validation, and test splits for each board type. These splits form the basis for all fine-tuning and evaluation setups described below. All models are fine-tuned using the Unsloth  [^2] https://unsloth.ai/ framework with a 4-bit quantization setup to reduce memory overhead. We apply parameter-efficient fine-tuning using LoRA with the following configuration: rank $r=16$, $\alpha=16$, dropout=$0.0$, batch size = $2$, learning rate = $2\mathrm{e}{-4}$, and the adam-8bit optimizer. This configuration was selected based on preliminary ablation experiments that balanced good trade-off between performance and training efficiency. Fine-tuning is performed for a single epoch, and two gradient accumulation steps, across all experiments. We use the same prompt template format described in Appendix A.1, ensuring consistency between training and evaluation. All experiments were conducted on a single NVIDIA A100 80GB GPU. To analyze generalization across board complexities, we conduct two focused fine-tuning setups that are discussed in the following sub-sections.
A.2.1 Simple-only Fine-Tuning
Models are trained exclusively on Simple Board (SB) data and evaluated on both SB and Regular Board (RB) test sets. The results are reported in Table 3 and Table 7. Given the similarity between the training and evaluation set for synthetic instructions (both drawn from low-ambiguity, structured templates), majority of the models achieve near-perfect Execution Success (score: 1.0), indicating strong memorization and overfitting.
Dataset SB
Training Set 1072
Validation Set 130
Test Set 130
RB
1168
130
130
TB
96
96
Table 2: Dataset statistics showing the number of boards in Simple Boards (SB), Regular Boards (RB), and number of scenarios in TidyBot (TB) across training, validation, and test splits.
Despite no exposure to human-written instructions during training, these models show competitive generalization to human-authored instructions on Simple boards. In several cases (e.g., LLaMA-3.3-70B and Qwen2.5-Coder-32B), performance surpasses that of closed-source models like Claude-3 and GPT-4. This suggests that structured synthetic pretraining on well-aligned prompts can transfer well to natural language instructions, in low-complexity environments.
While models perform strongly on the SB test split (comparable to full-dataset fine-tuning), their performance drops significantly on RB structures. This indicates limited generalization from simple to complex structures. However, the models exhibit reduced error rates (see Table 3) for both types of boards. This implies that the fine-tuned models are able to parse and respond to instructions with structurally valid outputs, even if the generated code is not matching with gold configuration. In other words, instruction-following behavior improves despite failure in task execution, highlighting partial generalization from simple to complex settings.
Error Rate
Synthetic Instructions Human-Authored Instructions
Simple Regular Simple Regular
Model Before After ($\downarrow$) Before After ($\downarrow$) Before After $\downarrow$) Before After ($\downarrow$)
CodeLlama-7B 0.23 0.00 0.79 0.00 0.45 0.00 0.74 0.00
CodeLlama-34B 0.61 0.00 0.82 0.00 0.79 0.00 0.72 0.00
Qwen2.5-Coder-7B 0.00 0.00 1.00 0.43 0.00 0.00 1.00 0.09
Qwen2.5-Coder-32B 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Qwen3-8B 0.90 0,00 0.98 0.98 1.00 0.00 0.95 0.99
Qwen3-32B 0.83 0.00 0.95 0.00 0.95 0.00 0.97 0.00
Llama-3.3-70B 0.00 0.00 1.00 0.00 0.00 0.00 0.72 0.00
Claude-3 0.00 - 0.00 - 0.00 - -
GPT-4o 0.00 - 0.00 - 0.00 - 0.28 -
Table 3: Error Rate results for models fine-tuned only on Simple Board (SB) instructions. We report performance before and after fine-tuning on both Synthetic and Human-Authored instructions across Simple and Regular board environments. Lower values indicate better adherence to the expected response format.
A.2.2 Regular-only Fine-Tuning
Models are trained exclusively on Regular Board (RB) data and evaluated on both Regular Board and Simple Board (SB) test sets. The results are summarized in Table 6 and Table 8. Fine-tuning on RB yields strong performance on the RB test split, with most models achieving high Execution Success rates (around 0.6) and consistently low Abort Rates (0.0), demonstrating effective instruction-following under high referential ambiguity and structural complexity.
Models also generalize better to SB than in the reverse setup (i.e., training on SB and testing on RB). While Execution Success on SB synthetic instructions remains modest overall, many models still produce valid, non-aborting responses, as reflected in low Abort Rates. This suggests that exposure to structurally complex instructions during training leads to more robust output formatting and moderate generalization to simpler scenarios.
That said, human instruction performance remains lower than synthetic in both RB and SB domains, highlighting challenges in bridging the synthetic–natural instruction gap—in less templated language.
A.2.3 Fine-Tuning on Combined Dataset
Models are fine-tuned jointly on the entire dataset comprising both Simple Board (SB) and Regular Board (RB) instructions. This setting provides balanced exposure to both low-ambiguity and high-ambiguity environments, allowing models to learn grounded execution strategies across a wider range of spatial reasoning tasks.
As discussed in the Section 5, models trained on the combined dataset achieve better performance on Simple boards test cases and moderate performance on Regular boards test set. The error rates (see Table 5) are reduced across all instruction styles and board types. This setting also results in the best overall generalization to human-authored instructions, indicating that training across diverse structural patterns is key to robust instruction-following behavior.
Score Ranges BLEU ES
SB RB SB RB
Median Values 0.356 0.024 0.979 0.623
Min-Max Range 0.356 0.024 0.979 0.623
Table 4: Median and min–max values of instruction similarity (BLEU) and Embedding Similarity (ES) across Simple Boards (SB) and Regular Boards (RB). Lower ES scores on RB highlight increased linguistic variability in human instructions for complex tasks.
Error Rate
Synthetic Instructions Human-Authored Instructions
Simple Regular Simple Regular
Model Before After ($\downarrow$) Before After ($\downarrow$) Before After $\downarrow$) Before After ($\downarrow$)
CodeLlama-7B 0.23 0.00 0.79 0.00 0.45 0.00 0.74 0.00
CodeLlama-34B 0.61 0.00 0.82 0.00 0.79 0.00 0.72 0.00
Qwen2.5-Coder-7B 0.00 0.00 1.00 0.00 0.00 0.00 1.00 0.00
Qwen2.5-Coder-32B 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Qwen3-8B 0.90 0,00 0.98 0.00 1.00 0.00 0.95 0.00
Qwen3-32B 0.83 0.00 0.95 0.00 0.95 0.00 0.97 0.00
Llama-3.3-70B 0.00 0.00 1.00 0.00 0.00 0.00 0.72 0.00
Claude-3 0.00 - 0.00 - 0.00 - 0.37 -
GPT-4o 0.00 - 0.00 - 0.00 - 0.28 -
Table 5: Error Rate results for models fine-tuned on both Simple Board (SB) and Regular Board instructions. We report performance before and after fine-tuning on both Synthetic and Human-Authored instructions across Simple and Regular board environments. Lower values indicate better adherence to the expected response format.
Error Rate
Synthetic Instructions Human-Authored Instructions
Simple Regular Simple Regular
Model Before After ($\downarrow$) Before After ($\downarrow$ Before After ($\downarrow$ Before After ($\downarrow$
CodeLlama-7B 0.23 0.00 0.79 0.00 0.45 0.00 0.74 0.00
CodeLlama-34B 0.61 0.00 0.82 0.00 0.79 0.00 0.72 0.00
Qwen2.5-Coder-7B 0.00 0.44 1.00 0.00 0.00 0.91 1.00 0.00
Qwen2.5-Coder-32B 0.00 0.99 0.00 0.00 0.00 1.00 0.00 0.62
Qwen3-8B 0.90 0.08 0.98 0.00 1.00 0.02 0.95 0.00
Qwen3-32B 0.83 0.08 0.95 0.02 0.95 0.05 0.97 0.07
Llama-3.3-70B 0.00 0.00 1.00 0.00 0.00 0.00 0.72 0.00
Claude-3 0.00 - 0.00 - 0.00 - 0.37
GPT-4o 0.00 - 0.00 - 0.00 - 0.28 -
Table 6: Error Rate results for models fine-tuned only on Regular Board (RB) instructions. We report performance before and after fine-tuning on both Synthetic and Human-Authored instructions across Simple and Regular board environments. Lower values indicate better adherence to the expected response format.
Execution Success
Synthetic Instructions Human-Authored Instructions
Simple Regular Simple Regular
Model Before After ($\uparrow$) Before After ($\uparrow$) Before After ($\uparrow$) Before After ($\uparrow$)
CodeLlama-7B 0.22 1.00 0.02 0.03 0.09 0.19 0.02 0.03
CodeLlama-34B 0.22 0.99 0.18 0.00 0.10 0.17 0.02 0.12
Qwen2.5-Coder-7B 0.53 1.00 0.00 0.01 0.07 0.45 0.00 0.00
Qwen2.5-Coder-32B 0.45 1.00 0.88 0.05 0.14 0.61 0.20 0.15
Qwen3-8B 0.00 1.00 0.00 0.00 0.00 0.45 0.00 0.00
Qwen3-32B 0.00 0.98 0.00 0.17 0.00 0.39 0.00 0.04
Llama-3.3-70B 0.34 1.00 0.00 0.35 0.11 0.69 0.07 0.18
Claude-3 0.87 - 0.86 - 0.43 - 0.25 -
GPT-4o 0.56 - 1.00 - 0.17 - 0.31 -
Table 7: Execution Success results for models fine-tuned only on Simple Board (SB) instructions. We report performance before and after fine-tuning on both Synthetic and Human-Authored instructions across Simple and Regular board environments. Higher values indicate better task success.
Execution Success
Synthetic Instructions Human-Authored Instructions
Simple Regular Simple Regular
Model Before After ($\uparrow$) Before After ($\uparrow$) Before After ($\uparrow$) Before After ($\uparrow$)
CodeLlama-7B 0.22 0.00 0.02 0.62 0.09 0.00 0.02 0.10
CodeLlama-34B 0.22 0.22 0.18 0.64 0.10 0.04 0.02 0.11
Qwen2.5-Coder-7B 0.53 0.51 0.00 0.61 0.07 0.04 0.00 0.23
Qwen2.5-Coder-32B 0.45 0.01 0.88 0.61 0.14 0.00 0.20 0.18
Qwen3-8B 0.00 0.24 0.00 0.61 0.00 0.15 0.00 0.15
Qwen3-32B 0.00 0.45 0.00 0.62 0.00 0.45 0.00 0.19
Llama-3.3-70B 0.34 0.13 0.00 0.59 0.11 0.17 0.07 0.13
Claude-3 0.87 - 0.86 - 0.43 - 0.25 -
GPT-4o 0.56 - 1.00 - 0.17 - 0.31 -
Table 8: Execution Success results for models fine-tuned only on Regular Board (RB) instructions. We report performance before and after fine-tuning on both Synthetic and Human-Authored instructions across Simple and Regular board environments. Higher values indicate better task success.
A.3 Instruction Similarity vs. Execution Success
We compute similarity between synthetic and human instructions associated with a given board configurations using two measures: BLEU (surface-level overlap) and Embedding cosine similarity using Sentence-BERT (semantic-level alignment).
Table 4 shows median and range values for both simple and regular board setups: (a) Instructions for simple boards show high semantic similarity between synthetic and human variants, leading to higher execution success; (b) Regular boards, which contain repeated structures, exhibit lower similarity, and correspondingly lower performance across models; (c) This supports our hypothesis that referential ambiguity, introduced by repetitive or structurally dense environments, reduces the efficacy of synthetic-only fine-tuning.
Figure 4: Prompt used for TidyBot evaluation. The upper block shows a sample instruction and its corresponding action sequence, while the lower block illustrates the input structure provided to the model. The task involves mapping natural language summaries to pick-and-place commands.
Figure 5: Example human-written instructions from the TidyBot dataset. These summaries describe high-level spatial organization tasks using natural, varied language.
A.4 TidyBot Dataset Evaluation
To assess real-world generalization, we evaluate models on the TidyBot dataset—an instruction-to-action benchmark for object arrangement in home environments Wu et al. (2023). TidyBot contains human-authored instructions describing semantic spatial tasks, such as placing utensils in drawers or aligning objects on tables and abstract spatial concepts, making it a suitable testbed for cross-domain transfer.
We perform few-shot evaluation on TidyBot using models fine-tuned solely on our synthetic data (SB or RB). We measure accuracy on seen vs. unseen objects and compare performance before and after fine-tuning to assess transferability and Seen vs. Unseen Generalization. These results (see Table 1) highlight the best performing models (Llama3.3-70B and Qwen2.5-32B) are able to generalize well on a new dataset.
Figure 6: Web interface used for human reconstruction of spatial configurations. Instructions appear on the left, while the right panel allows annotators to place objects on an 8×8 grid using a visual palette. The interface supports actions such as object selection and cell duplication to aid in building complex or repetitive structures.
A.5 Human Evaluation
To establish a human performance baseline, we developed an interactive web interface (Figure 6) that displays the instruction on the left and an editable 8$\times$8 grid on the right. Annotators were tasked with reconstructing the spatial configuration described by the input instruction using the grid. The interface included features such as cell copy-paste to efficiently handle repeated structures and embedded step-by-step guidelines to standardize the reconstruction process.
An annotator was recruited to perform the task. While the annotator had no prior experience in spatial layout tasks, they were familiar with general data labeling workflows. The evaluation involved reconstructing a total of 260 boards (130 Simple and 130 Regular) and required approximately 30 hours to complete.
The reconstructed boards were programmatically (by executing the constructed scripts as a result of interaction) compared to the corresponding gold configurations. The resulting match scores are reported as human baseline accuracies and serve as an upper bound reference for model performance.
template A.5.1
System Info
You are a helpful assistant who is designed to interpret and translate natural language instructions into python executable code snippets.
Environment Info
The environment is an 8x8 grid allowing shape placement and stacking. A shape can be placed in any cell, while stacking involves adding multiple shapes to the same cell, increasing its depth. Shapes typically occupy a single cell, except for the "bridge," which spans two cells and requires two other shapes for stacking. Horizontal bridges span adjacent columns (left and right), and vertical ones span consecutive rows (top and bottom). Stacking is only possible if the shapes have matching depths.
In the grid, columns align with the x-axis and rows with the y-axis. Python indexing is used to identify each cell. The cell in the top-left corner is in the first row and first column, corresponding to x and y values of 0, 0. Similarly, the top-right corner cell is in the first row and eighth column, with x and y values of 0, 7.
- Use the shape name ’bridge-h’ if a bridge is placed horizontally
- Use the shape name ’bridge-v’ if a bridge is placed vertically
The following functions are already defined; therefore, do not generate additional code for it
- Use ‘put(board: np.ndarray, shape: string, color: string, x: int, y: int) to place a shape on the board
Task Info
For each instruction labeled Instruction: please respond with code under the label Function: followed by a newline and usage for the function under the label Usage: followed by a newline.
Context Info
\$INCONTEXT_SAMPLES
Other Info
Do not generate any other text/explanations.
The order of colors, x, y matters, as these are assigned to the shapes in the same sequence. Ensure the response can be executed by Python ‘exec()‘, e.g.: no trailing commas, no periods, etc.
Let’s begin
Instruction:
\$TEST_INSTRUCTION
Figure 7: Prompt template used for the spatial grounding task for Simple boards. The system information specifies system level behavior, the environment information indicates the environment details of the user-agent environment, the context information describes the in-context examples, task information indicates the specific response format to follow.
template A.5.2
System Info
You are a helpful assistant who is designed to interpret and translate natural language instructions into python executable code snippets.
Environment Info
The environment is an 8x8 grid allowing shape placement and stacking. A shape can be placed in any cell, while stacking involves adding multiple shapes to the same cell, increasing its depth. Shapes typically occupy a single cell, except for the "bridge," which spans two cells and requires two other shapes for stacking. Horizontal bridges span adjacent columns (left and right), and vertical ones span consecutive rows (top and bottom). Stacking is only possible if the shapes have matching depths.
In the grid, columns align with the x-axis and rows with the y-axis. Python indexing is used to identify each cell. The cell in the top-left corner is in the first row and first column, corresponding to x and y values of 0, 0. Similarly, the top-right corner cell is in the first row and eighth column, with x and y values of 0, 7.
- Use the shape name ’bridge-h’ if a bridge is placed horizontally - Use the shape name ’bridge-v’ if a bridge is placed vertically
The following functions are already defined; therefore, do not generate additional code for it
- Use ‘put(board: np.ndarray, shape: string, color: string, x: int, y: int)‘ to place a shape on the board
- Use ‘\$COMBO_NAME(board: np.ndarray, colors: \$COLORS, x: int, y: int)‘ to place a ’\$COMBO_NAME’ object on the board
Task Info
For each instruction labeled Instruction: please respond with code under the label Output: followed by a newline.
Context Info
\$INCONTEXT_SAMPLES
Other Info
Do not generate any other text/explanations.
The order of colors, x, y matters, as these are assigned to the shapes in the same sequence. Ensure the response can be executed by Python ‘exec()‘, e.g.: no trailing commas, no periods, etc.
Lets begin
Instruction:
\$TEST_INSTRUCTION
Figure 8: Prompt template used for the spatial grounding task for Regular boards. The system information specifies system level behavior, the environment information indicates the environment details of the user-agent environment, the context information describes the in-context examples, task information indicates the specific response format to follow.